OpenStudy (tootsi123):

The sum of two numbers is 55 . One number is 4 times as large as the other. What are the numbers?

OpenStudy (tootsi123):

I have tried to love this on my own and this is how i did it 55/2= 27.5 and then i divided that by 4. I know i did this wrong so please explain

OpenStudy (anonymous):

one number is n, the other one is 4n and n + 4n = 55

OpenStudy (tootsi123):

Okay... so i divide 55 by 5 then right? i am saying this because 4n+n = 5n

OpenStudy (anonymous):

yes divide by 5

OpenStudy (jjwalton04):

This is what i got x + 4x = 55 5x = 55 x = 11 4x = 44

OpenStudy (jjwalton04):

Therefor the numbers are 11 and 44

OpenStudy (tootsi123):

:) got it

OpenStudy (jjwalton04):

Srry i have a headache but what did u get?

OpenStudy (tootsi123):

Its okay.. i have a headache to part of the reason i am having trouble concentrating.... and what you said about before about what you got

OpenStudy (jjwalton04):

OK thats good lol

OpenStudy (jjwalton04):

well not the part with the headache lolz

OpenStudy (tootsi123):

Yeah i know :)

OpenStudy (jjwalton04):

Well if u ever need my help again, just tag me

OpenStudy (tootsi123):

Okay well i need help with this simplify (x^4)^2 without parenthesis

OpenStudy (jjwalton04):

ok hld on

OpenStudy (tootsi123):

I figured it out... its x^8 but thank you anyway

OpenStudy (jjwalton04):

oh lolz

OpenStudy (jjwalton04):

u sure?

OpenStudy (jjwalton04):

Bc i got x^16

OpenStudy (tootsi123):

how?? I did 4*2 which is eight but wait i think i am forgetting that you have to multiply that by two also because it is on the outside of the parenthesis and i right

OpenStudy (jjwalton04):

Exactly.

OpenStudy (tootsi123):

Oh okay thanks